Solve the following problem and answer what is being asked. Express all answers in four decimal places.

A piece of metal was heated and placed in a calorimeter containing ice. Assuming no heat escapes the calorimeter, solve what is being asked given the following data:

Initial Temperature of metal: 98
1
Expert's answer
2021-10-01T12:44:02-0400


Heat loss by the metal:

"Q_m=mcΔT \\

= 42.4 \times 0.38 \times (1.6-98) \\

= -1553.2 \;J"

Experimental heat of fusion:

"Q_E = \frac{Q_m + m_{ice} cΔT}{-m_{ice}} \\

= \frac{-1553.2 +4.6 \times 2.04 \times 6.4}{-4.6} \\

= +324.6 \;J"

% error "= |\frac{Q_E-Q_T}{Q_T}| \times 100"

"= |\frac{324.6-334}{334}| \times 100 \\

= 2.81 \; \%"